### v1

@@ 4.4 Detection probability in the SFR- plane in intensity
In the previous sections, we characterized the classical confusion limit only in term of flux density. However, to understand its impact on the observatory science, it is essential to consider the impact on intrinsic physical properties. We thus computed the probability to detect a galaxy above the classical confusion limit as a function of SFR and redshift. The border between the regions of low and high probability is blurred, since our model has a diversity of SEDs and our simulation produces a completeness curve which have a continuous transition from 0 to 1 (see Sect. 4.2). The results are presented in Fig. 7 together with the tracks corresponding to galaxies of various stellar masses following the main sequence relation of Schreiber et al. (2015) and the evolution of the knee of the infrared luminosity function (L, Traina et al., 2024).
For the shortest wavelength (PHI1_1 band centered on 25 m, upper left corner of Fig. 7), the 10 M and 10 M main-sequence galaxies are recovered up to 2.5 and 3.5, respectively. The L galaxies are slightly above the detection border up to z3 and undetected above. Overall, the border between detections and non detections moves towards higher SFR with increasing redshift. However, we can identify some specific features. At 2, we are probing the 10 m rest-frame dip in SED between the various PAH bands, and galaxies are harder to detect (higher SFR limit). At 2, the 7.7 m PAH band enters the representative filters making the galaxies easier to detect (lower SFR limit). Around , the typical SFR at which galaxies are detected increases sharply, and only some rare outliers can be detected. This is the consequence of the absence of strong dusty features below 6 m rest-frame in the SED of star-forming galaxies.
@@
Finally, we combined all the bands to derive the probability to detect galaxies in at least one band (lower right corner of Fig. 7). This illustrates the parameter space, which could be probed by PRIMAger above the classical confusion limit. A galaxy exactly on the main-sequence relation and with a stellar mass of 10 M and 10 M can be detected up to 2.5, and 5, respectively. The L galaxies are detected up to z3.5. The border between detection and non detection is almost featureless, since the PAH slides through the various representative sub-filters. This illustrates how hyperspectral imaging can help to deal with confusion. However, the dip at 1.5 seen in PHI1_1 is still present, since there is no shorter wavelength to observe this redshift range around 7.7 m rest-frame.
@@ 5 Confusion in polarization
In this section, we discuss the impact of confusion on polarization data.
@@ 5.1 Photometric accuracy
In Fig. 8 (left panels), we show the ratio between the brightest galaxy polarized flux density in the beam and the measured value in the simulated map. In contrast with the intensity maps (Sect. 4.1 and Fig. 3, we do not observe any bias in the median flux ratio. This is likely to be because the contribution of several sources in the beam is not fully additive if their polarization angles are not aligned. However, similarly to the intensity, we still observe an increase of the half width of the 1 confidence region from 9 % to 20 % from PPI1 to PPI4, but overall the dispersion is slightly lower than it is for intensity.
We thus recover the polarized flux density of sources just above the classical confusion limit with a good accuracy, while it is not the case in intensity (see Sect. 4.1). However, the polarized flux density is much weaker than the intensity, and detecting it will requires much deeper data. In addition, as discussed in Sect. 2.3, we did not include galaxy alignments, which could produce a small flux density excess similarly to what happens in intensity.
@@ 5.2 Recovering polarized angles and polarized fraction
The polarized fraction () is another useful quantity to characterize distant unresolved galaxies. We derived for each galaxy detected in the map, extracting the value of at the same position in the intensity map. In Fig. 8 (central panels), we show the polarized fraction ratio between the brightest galaxy in the beam (highest ) and the measurement in the simulated map. At low polarized flux, the intrinsic polarized fraction of the brightest galaxy is significantly larger than the measured one. This is a natural consequence of the negligible bias found for the polarized flux density measurements (, Sect. 5.1) and the significant bias found in intensity (, Sect. 4.1 and Fig. 3), since with the first factor being close to one and the second being significantly above (i.e. approximately the inverse of the quantity shown in Fig. 3).
@@ 5.3 Purity and completeness
The purity of the samples extracted from polarization maps is excellent (98 %, see Fig. 6), and is barely effected by the choice of definition of the clustering. This is not surprising as the low level of flux boosting by the neighbors on means definitions A and B consider very similar matches. Finally, the clustering is also not expected to have a strong impact, since the polarized flux density does not add up as it does intensity data.
The completeness curves as a function of the intrinsic galaxy have a rather similar shape to those found for , but the transition between low and high completeness appears at lower flux densities (see Fig. 9). As shown in Table 2 summarizing the completeness in both intensity and polarization, the flux density limits in polarization is up to a factor of 1.8 lower than the product of the limit in intensity by the polarized fraction. This is again probably caused by the the non-additivity of the flux density inside a beam in polarization mitigating slightly the blending problems. Consequently, the surface density of sources above the classical confusion limit in a given band is higher in polarization than in intensity (see Table 2).
Finally, we find that the impact of clustering on the 50 % completeness flux density and the classical confusion limit is negligible (see Fig. 5). This is the consequence of the confusion in polarization being driven by chance polarization alignments rather than the local source density. There is a small impact (10 %) on the 80 % completeness flux density, which could have the same cause as the effect seen at short wavelength in intensity (see discussion in Sect. 4.2).
@@ 5.4 Polarized detection probability in the SFR- plane above the classical confusion limit
We also studied the probability of recovering a source at the classical confusion limit using the same method as described in Sect. 4.4 for the intensity. The results are shown in Fig. 10.
In the PPI bands, only z2.5 galaxies emerge from the confusion, similarly to intensity. The PPI1 band probes 10 M, 10 M, and L galaxies up to z0.5, z1.5, and z0.5, respectively. Above , the probability of detection remains small even for the most strongly star-forming galaxies. Although they are less sensitive at low redshift, the PPI2 and PPI3 bands are slightly better at catching these extreme sources, since they observe them closer to their peak of emission. Consequently, the probability to detect a source in at least one band is very similar to the probability to detect it in the PPI1 band with the exception of the the tail at z2.25 and SFR1000 M/yr.
@@
| CIB 1 fluctuations at native resolution in kJy/sr | 1.7 | 2.6 | 3.0 | 2.4 |
@@
| CIB 1 fluctuations at PPI4 angular resolution in kJy/sr | 1.2 | 1.8 | 2.3 | 2.4 |
@@
| Surface brightness limit to obtain a 10 % uncertainty on the color with PPI4 in kJy/sr | 11 | 19 | 20 | – |
| Correlation coefficient between a band and PPI4 at the surface brightness limit | 0.79 | 0.87 | 0.96 | – |
@@
The confusion noise is not only a problem for studying high-redshift galaxies. The fluctuations of the polarized CIB can impact both diffuse foreground and background measurements. The case of the cosmic microwave background has already been extensively discussed by Lagache et al. (2020). In this section, we evaluate the impact on diffuse Galactic emission and nearby, spatially-resolved, galaxies.
@@
The fluctuations measured in presence of a strong foreground are up to 20 % higher than in the CIB-only case. Hence, this is a small but non-negligible effect. If we had masked only the positive 5 outliers, the CIB fluctuations would have been up to 50 % higher in the strong foreground case, but unchanged in the pure CIB case since there are no strong negative fluctuations. Our table also shows that CIB fluctuations at PPI4 resolution are lower than at native resolution. In surface brightness units, the signal from the constant foreground does not vary with the beam size, while a larger beam contains more sources and reduces the stochastic fluctuations.
Finally, we explored the impact of CIB on foreground color measurements. We use the “astrodust” model of dust emission and polarization (Hensley & Draine, 2023) and have assumed that the dust is heated by a radiation field appropriate for diffuse atomic gas, to derive nominal input values of the PPI1/PPI4, PPI2/PPI4, and PPI3/PPI4 foreground colors of 0.52, 1.02, and 1.25, respectively. We varied the foreground surface brightness fixing the input color, and derived the relative uncertainty on the measured foreground color produced by CIB fluctuations. We then interpolated between these values to determine the surface brightness sensitivity limit corresponding to a 10 % uncertainty (see Table 3). The foreground surface brightness limits to reach a 10 % precision on the foreground color are lower than 10 times the 1 CIB fluctuations, which is the limit expected based only on the numerator part of the color computation. However, as shown in the last row of Table 3, the confusion noise is highly correlated. Positive and negative fluctuations of the CIB are thus expected to impact both bands in a similar way, reducing their impact on the ratio. The correlation between bands thus mitigates the confusion noise in such analyses.
| Band | P | N | Mean | Mean SFR |
| Jy | M/yr | |||
| Deep field (1500h, 1deg) | ||||
| with payload sensitivity | ||||
@@
| Wide field (1500h, 10deg) | ||||
| with payload sensitivity | ||||
@@
| Deep field (1500h, 1deg) | ||||
| with baseline sensitivity | ||||
@@
| Wide field (1500h, 10deg) | ||||
| with baseline sensitivity | ||||
@@ 6.1 Expected impact of confusion in intensity
In our analysis, we will consider two cases for the instrumental noise. The payload sensitivity is the guaranteed performance of the instrument. It is a very conservative estimate used to prove the feasibility of the PI science. The real performance is expected to be significantly better. We thus use a second estimate based on the baseline sensitivity model. To be consistent with the confusion, we also consider a 5 limit. We treat the case of two fields observed 1 500 h each. The deep field has a 1 deg area, while the wide field covers 10 deg.
In Fig. 12, we compare the classical confusion limit with the instrumental sensitivity. The confusion has a steeper rise with increasing wavelength than the sensitivity for both sensitivity estimates. Consequently, the short wavelengths are noise-limited and the long wavelengths are confusion-limited. For the wide survey and a payload sensitivity, the curves cross at 65 m around the center of band PHI2. As shown in Sect. 4.4 and Fig. 7, the red end of the PHI2 and the PPI bands at the confusion limit are not probing a part of the SFR- space missed by the other sub-bands. However, PPI data are important to characterize the physics of the objects, and PHI priors will be crucial to deblend them. In the deep field with a payload sensitivity, the classical confusion limit is reached at around 45 m, and the PHI2 band will thus be affected by it. At this wavelength, the 7.7 m PAH feature can be seen up to z5. (see Sect. 4.4). This ensures that we will benefit from very deep priors up to this redshift before reaching the confusion, which will be essential to deblend the PHI2 and PPI bands and obtain more accurate physical constraints.
For the baseline sensitivity, the confusion is reached at 55 m and 45 m in the wide and deep fields, respectively. We will thus be confusion limited in the PHI2 band, except in its bluest part for the wide fields. The SFR- space probed will thus be similar to the pure confusion case discussed in Sect. 4.4 and Fig. 7. We also note that in PPI bands, the instrumental noise will be 2.5 orders of magnitude below the confusion. These data will thus be extremely close to the noiseless case discussed in this paper and will be ideal to apply modern deep learning deblending algorithm (e.g., Lauritsen et al., 2021).
@@ 6.2 Feasibility of dust polarization surveys of distant galaxies
Far-infrared blank-field polarization surveys are in a totally uncharted territory. With our analysis, we can now set constraints on the expected classical confusion limit and we have demonstrated that we can recover constraints on the polarized flux density and angle of a galaxy (Sect. 5). However, since the signal will be fainter than in intensity, it is important to check if the instrumental sensitivity will be good enough to detect a large sample of sources.
In Fig. 12, we also compare the confusion and the instrumental limits in polarization. In PPI bands, the payload and the baseline sensitivities are very different. In the first case, the sensitivity limit in the deep fields are about an order of magnitude above the classical confusion limit (1.5 dex above for the wide). In the second more optimistic case, the wide field is noise limited. In the deep field, the PPI1 and PPI2 band are close to the classical confusion limit, while the other bands have a sensitivity limit slightly below the classical confusion limit. This means that we should be very close to the confusion limited case in the SFR- plane discussed in Sect. 7.
@@
For the payload sensitivity, we expect 100 galaxies per field in PPI1, but fewer than 10 in PPI4. Since it is a totally unexplored parameter space, it will open a new window with small but statistical significant samples. In the deep field, half of the sources detected in polarization are below , and only a handful are above with a tail up to . We will thus trace mainly intermediate redshifts, though the dust polarization properties of galaxies at these epochs are currently totally unexplored. In terms of SFR, we will span a large range of SFR from nearby 0.2 M/yr to high- 1000 M/yr galaxies. In the wide field, objects are detected only up to . As expected, the lowest SFR will not be probed. Paradoxically, we will also observe less 100 M/yr systems than in the deep field. This is driven by the very low number density of these high-SFR objects at z0.4 and the polarized flux density limit being too high to be able to detect any high- system.
With the baseline sensitivity, the confusion and instrumental noise will be similar in the PPI1 and PPI2 bands. Several thousands of galaxies will be detected both in the deep and wide fields (Table 4). The number of detections in the deep field is a factor of 2 smaller than in the wide field. The PPI1 band will provide the highest number of detections, but even the PPI4 band will detect several hundred sources allowing statistical studies of the polarized SEDs. Both deep and wide fields have a peak redshift distribution around z0.4 with large tail up to . In terms of SFR, we will cover 5 orders of magnitudes from 0.01 to 1000 M/yr.
The payload sensitivity would open a new window on the dust polarization of high-redshift galaxies with more than hundred detections up to z1.5. With the baseline sensitivity, the results would be totally transformational by opening this new window directly with several thousands of sources up to .

### v3

@@ 4.4 Detection probability in the SFR- plane in intensity
In the previous sections, we characterized the classical confusion limit only in term of flux density. However, to understand its impact on the observatory science, it is essential to consider the impact on intrinsic physical properties. We thus computed the probability to detect a galaxy above the classical confusion limit as a function of SFR and redshift. The border between the regions of low and high probability is blurred, since our model has a diversity of SEDs and our simulation produces a completeness curve which have a continuous transition from 0 to 1 (see Sect. 4.2). The results are presented in Fig. 7 together with the tracks corresponding to galaxies of various stellar masses following the main sequence relation of Schreiber et al. (2015) and the evolution of the knee of the infrared luminosity function (L⋆, Traina et al., 2024).
For the shortest wavelength (PHI1_1 band centered on 25 m, upper left corner of Fig. 7), the 1010 M⊙ and 1011 M⊙ main-sequence galaxies are recovered up to 2.5 and 3.5, respectively. The L⋆ galaxies are slightly above the detection border up to z3 and undetected above. Overall, the border between detections and non detections moves towards higher SFR with increasing redshift. However, we can identify some specific features. At 2, we are probing the 10 m rest-frame dip in SED between the various PAH bands, and galaxies are harder to detect (higher SFR limit). At 2, the 7.7 m PAH band enters the representative filters making the galaxies easier to detect (lower SFR limit). Around , the typical SFR at which galaxies are detected increases sharply, and only some rare outliers can be detected. This is the consequence of the absence of strong dusty features below 6 m rest-frame in the SED of star-forming galaxies.
@@
Finally, we combined all the bands to derive the probability to detect galaxies in at least one band (lower right corner of Fig. 7). This illustrates the parameter space, which could be probed by PRIMAger above the classical confusion limit. A galaxy exactly on the main-sequence relation and with a stellar mass of 1010 M⊙ and 1011 M⊙ can be detected up to 2.5, and 5, respectively. The L⋆ galaxies are detected up to z3.5. The border between detection and non detection is almost featureless, since the PAH slides through the various representative sub-filters. This illustrates how hyperspectral imaging can help to deal with confusion. However, the dip at 1.5 seen in PHI1_1 is still present, since there is no shorter wavelength to observe this redshift range around 7.7 m rest-frame.
@@ 5 Confusion in polarization
In this section, we discuss the impact of confusion on polarization data. We do not take into account the instrumental noise in this section. If we assume a different mean polarized fraction , both the confusion noise and the source polarized flux density scales as (/1 %). Consequently, the x-axis of Fig. 8 and 9 and the values in Table 3 must be shifted by this factor, while Fig. 10 is unchanged.
@@ 5.1 Photometric accuracy
In Fig. 8 (left panels), we show the ratio between the brightest galaxy polarized flux density in the beam and the measured value in the simulated map. In contrast with the intensity maps (Sect. 4.1 and Fig. 3), we do not observe any bias in the median flux ratio. This is likely to be because the contribution of several sources in the beam is not fully additive if their polarization angles are not aligned. However, similarly to the intensity, we still observe an increase of the half width of the 1 confidence region from 9 % to 20 % from PPI1 to PPI4, but overall the dispersion is slightly lower than it is for intensity.
We can thus recover the polarized flux density of sources just above the classical confusion limit with a good accuracy, while it is not the case in intensity (see Sect. 4.1). However, the polarized flux density is much weaker than the intensity, and detecting it will require much deeper data. In addition, as discussed in Sect. 2.3, we did not include galaxy alignments, which could produce a small polarized flux density excess similarly to what happens in intensity.
@@ 5.2 Recovering polarized angles and polarized fraction
The polarized fraction () is another useful quantity to characterize distant unresolved galaxies. We derived for each galaxy detected in the map, extracting the value of at the same position in the intensity map. In Fig. 8 (central panels), we show the polarized fraction ratio between the brightest galaxy in the beam (highest ) and the measurement in the simulated map. At low polarized flux, the intrinsic polarized fraction of the brightest galaxy is significantly larger than the measured one. This is a natural consequence of the negligible bias found for the polarized flux density measurements (, Sect. 5.1) and the significant bias found in intensity (, Sect. 4.1 and Fig. 3), since with the first factor being close to one and the second being significantly above (i.e., the inverse of the quantity shown in Fig. 3).
@@ 5.3 Purity and completeness
The purity of the samples extracted from polarization maps is excellent (98 %, see Fig. 6), and is barely affected by the choice of definition of the clustering. This is not surprising as the low level of flux boosting by the neighbors on means definitions A and B consider very similar matches. Finally, the clustering is also not expected to have a strong impact, since the polarized flux density does not add up as it does for the intensity data.
The completeness curves as a function of the intrinsic galaxy have a rather similar shape to those found for , but the transition between low and high completeness appears at lower flux densities (see Fig. 9). As shown in Table 2 summarizing the completeness in both intensity and polarization, the polarized flux density limits are up to a factor of 1.8 lower than the product of the limits in intensity by the mean polarization fraction . This is again probably caused by the the non-additivity of the polarized flux density inside a beam in polarization mitigating slightly the blending problems. Consequently, the surface density of sources above the classical confusion limit in a given band is higher in polarization than in intensity (see Table 2).
Finally, we find that the impact of clustering on the 50 % completeness polarized flux density and the classical confusion limit is negligible (see Fig. 5, red lines). This is the consequence of the confusion in polarization being driven by chance polarization alignments rather than the local source density. There is a small impact (10 %) on the 80 % completeness polarized flux density, which could have the same cause as the effect seen at short wavelength in intensity (see discussion in Sect. 4.2).
@@ 5.4 Polarized detection probability in the SFR- plane above the classical confusion limit
We also studied the probability of recovering a source above the classical confusion limit using the same method as described in Sect. 4.4 for the intensity. The results are shown in Fig. 10.
In the PPI bands, only z2.5 galaxies emerge from the confusion, similarly to intensity. The PPI1 band probes 1010 M⊙, 1011 M⊙, and L⋆ galaxies up to z0.5, z1.5, and z0.5, respectively. Above , the probability of detection remains small even for the most strongly star-forming galaxies. Although they are less sensitive at low redshift, the PPI2 and PPI3 bands are slightly better at catching these extreme sources, since they observe them closer to their peak of emission. Consequently, the probability to detect a source in at least one band is very similar to the probability to detect it in the PPI1 band with the exception of the the tail at z2.25 and SFR1000 M⊙/yr.
@@
| Polarized CIB 1 fluctuations at native resolution in kJy/sr | 1.7 | 2.6 | 3.0 | 2.4 |
@@
| Polarized CIB 1 fluctuations at PPI4 angular resolution in kJy/sr | 1.2 | 1.8 | 2.3 | 2.4 |
@@
| Polarized surface brightness limit to obtain a 10 % uncertainty on the color with PPI4 in kJy/sr | 11 | 19 | 20 | – |
| Correlation coefficient between a band and PPI4 at the polarized surface brightness limit | 0.79 | 0.87 | 0.96 | – |
@@
The confusion noise is not only a problem for studying high-redshift galaxies. The fluctuations of the polarized CIB can impact both diffuse foreground and background measurements. The case of the cosmic microwave background has already been extensively discussed by Lagache et al. (2020). While the contribution of astrophysical components in intensity is additive, the polarization is a vectorial quantity and leads to a more complex combination of the various components. In intensity, the confusion noise from the CIB only is sufficient to estimate its impact on the foreground science (e.g., diffuse Galactic emission and nearby or spatially-resolved galaxies). In contrast, as we will show in this section, the impact of the CIB in polarization depends on the foreground polarized surface brightness.
@@
The fluctuations measured in presence of a strong foreground are up to 20 % higher than in the CIB-only case. Hence, this is a small but non-negligible effect. If we had masked only the positive 5 outliers, the CIB fluctuations would have been up to 50 % higher in the strong foreground case, but unchanged in the pure CIB case since there are no strong negative fluctuations. Our table also shows that CIB fluctuations at PPI4 resolution are lower than at native resolution. In polarized surface brightness units, the signal from the constant foreground does not vary with the beam size, while a larger beam contains more sources and reduces the stochastic fluctuations.
Finally, we explored the impact of CIB on foreground color measurements. We use the “astrodust” model of dust emission and polarization (Hensley & Draine, 2023) and have assumed that the dust is heated by a radiation field appropriate for diffuse atomic gas, to derive nominal input values of the PPI1/PPI4, PPI2/PPI4, and PPI3/PPI4 foreground colors of 0.52, 1.02, and 1.25, respectively. We varied the foreground polarized surface brightness fixing the input color, and derived the relative uncertainty on the measured foreground color produced by CIB fluctuations. We then interpolated between these values to determine the polarized surface brightness sensitivity limit corresponding to a 10 % uncertainty (see Table 3). The foreground polarized surface brightness limits to reach a 10 % precision on the foreground color are lower than 10 times the 1 CIB fluctuations, which is the limit expected based only on the numerator part of the color computation. However, as shown in the last row of Table 3, the confusion noise is highly correlated. Positive and negative fluctuations of the CIB are thus expected to impact both bands in a similar way, reducing their impact on the ratio. The correlation between bands thus mitigates the confusion noise in such analyses.
We thus showed that the confusion noise from the polarized CIB depends on the properties of the foreground and is also strongly correlated between bands. Our work provides first estimates of the impact of CIB to study Galactic emission and nearby galaxies in polarization. More complex simulations including full foreground models will be key to prepare these science cases.
| Band | Plim | Ndet | Mean | Mean SFR |
| Jy | M⊙/yr | |||
| Deep field (1500h, 1deg2) | ||||
| with payload required sensitivity | ||||
@@
| Wide field (1500h, 10deg2) | ||||
| with payload required sensitivity | ||||
@@
| Deep field (1500h, 1deg2) | ||||
| with conservative estimated sensitivity | ||||
@@
| Wide field (1500h, 10deg2) | ||||
| with conservative estimated sensitivity | ||||
@@ 6.1 Expected impact of confusion in intensity
In our analysis, we will consider two cases for the instrumental noise. The payload required sensitivity is the guaranteed performance of the instrument. It is a very conservative estimate to ensure high confidence in meeting the PI science goals. The real performance is expected to be much better, with at least 60 % margin and in some cases by factors of several. We consider an intermediate estimate termed the conservative estimated sensitivity which lies between the payload requirement and the actual estimated performance. To be consistent with the confusion, we also consider a 5 limit. We treat the case of two fields observed 1 500 h each. The deep field has a 1 deg2 area, while the wide field covers 10 deg2.
In Fig. 12, we compare the classical confusion limit with the instrumental sensitivity. The confusion has a steeper rise with increasing wavelength than the sensitivity for both sensitivity estimates. Consequently, the short wavelengths are noise-limited and the long wavelengths are confusion-limited. For the wide survey and a payload required sensitivity, the curves cross at 65 m around the center of band PHI2. As shown in Sect. 4.4 and Fig. 7, the red end of the PHI2 and the PPI bands at the confusion limit are not probing a part of the SFR- space missed by the other sub-bands. However, PPI data are important to characterize the physics of the objects, and PHI priors will be crucial to deblend them. In the deep field with a payload required sensitivity, the classical confusion limit is reached at around 45 m, and the PHI2 band will thus be affected by it. At this wavelength, the 7.7 m PAH feature can be seen up to z5. (see Sect. 4.4). This ensures that we will benefit from very deep priors up to this redshift before reaching the confusion, which will be essential to deblend the PHI2 and PPI bands and obtain more accurate physical constraints.
For the conservative estimated sensitivity, the confusion is reached at 55 m and 45 m in the wide and deep fields, respectively. We will thus be confusion limited in the PHI2 band, except in its bluest part for the wide fields. The SFR- space probed will thus be similar to the pure confusion case discussed in Sect. 4.4 and Fig. 7. We also note that in PPI bands, the instrumental noise will be 2.5 orders of magnitude below the confusion. These data will thus be extremely close to the noiseless case discussed in this paper and will be ideal to apply modern deep learning deblending algorithm (e.g., Lauritsen et al., 2021).
@@ 6.2 Feasibility of dust polarization surveys of distant galaxies
Far-infrared blank-field polarization surveys are in a totally uncharted territory. With our analysis, we can now set constraints on the expected classical confusion limit and we have demonstrated that we can recover constraints on the polarized flux density and angle of a galaxy (Sect. 5). However, since the signal will be fainter than in intensity, it is important to check if the instrumental sensitivity will be good enough to detect a large sample of sources. In this section, we discuss only our standard simulation assuming a mean polarization fraction of 1 %. In Appendix C, we consider the alternative cases with 3-lower (=0.7 %) and 3-higher (=1.3 %) values (see Sect. 2.3).
In Fig. 12, we also compare the confusion and the instrumental limits in polarization. In PPI bands, the payload required and the conservative estimated sensitivities are very different. In the first case, the sensitivity limit in the deep fields are about an order of magnitude above the classical confusion limit (1.5 dex above for the wide). In the second more optimistic case, the wide field is noise limited. In the deep field, the PPI1 and PPI2 band are close to the classical confusion limit, while the other bands have a sensitivity limit slightly below the classical confusion limit. This means that we should be very close to the confusion limited case in the SFR- plane discussed in Sect. 7.
@@
For the payload required sensitivity, we expect 100 galaxies per field in PPI1, but fewer than 10 in PPI4. Since it is a totally unexplored parameter space, it will open a new window with small but statistical significant samples. In the deep field, half of the sources detected in polarization are below , and only a handful are above with a tail up to . We will thus trace mainly intermediate redshifts, though the dust polarization properties of galaxies at these epochs are currently totally unexplored. In terms of SFR, we will span a large range of SFR from nearby 0.2 M⊙/yr to high- 1000 M⊙/yr galaxies. In the wide field, objects are detected only up to . As expected, the lowest SFR will not be probed. Paradoxically, we will also observe less 100 M⊙/yr systems than in the deep field. This is driven by the very low number density of these high-SFR objects at z0.4 and the polarized flux density limit being too high to be able to detect any high- system.
With the conservative estimated sensitivity, the confusion and instrumental noise will be similar in the PPI1 and PPI2 bands. Several thousands of galaxies will be detected both in the deep and wide fields (Table 4). The number of detections in the deep field is a factor of 2 smaller than in the wide field. The PPI1 band will provide the highest number of detections, but even the PPI4 band will detect several hundred sources allowing statistical studies of the polarized SEDs. Both deep and wide fields have a peak redshift distribution around z0.4 with large tail up to . In terms of SFR, we will cover 5 orders of magnitudes from 0.01 to 1000 M⊙/yr.
The payload required sensitivity would open a new window on the dust polarization of high-redshift galaxies with more than hundred detections up to z1.5. With the conservative estimated sensitivity, the results would be totally transformational by opening this new window directly with several thousands of sources up to .